# MROSS: Multi-Round Region-Based Optimization for Scene Sketching

### v1

@@ Abstract
Scene sketching is to convert a scene into a simplified, abstract representation that captures the essential elements and composition of the original scene. It requires semantic understanding of the scene and consideration of different regions within the scene. Since scenes often contain diverse visual information across various regions, such as foreground objects, background elements, and spatial divisions, dealing with these different regions poses unique difficulties. In this paper, we define a sketch as some sets of Bézier curves. We optimize the different regions of input scene in multiple rounds. In each round of optimization, strokes sampled from the next region can seamlessly be integrated into the sketch generated in the previous round of optimization. We propose additional stroke initialization method to ensure the integrity of the scene and the convergence of optimization. A novel CLIP-Based Semantic loss and a VGG-Based Feature loss are utilized to guide our multi-round optimization. Extensive experimental results on the quality and quantity of the generated sketches confirm the effectiveness of our method.
Figure 1: Our method converts a scene image into a sketch with different levels of abstraction. The first row presents a concise, comprehensive vector sketch of the scene, ideal for designer edits. The second row illustrates the multi-round optimization process, achieving coarse-to-concrete sketches without altering the total stroke count.
Introduction
@@
Scene sketching offers the advantage of conveying information quickly and efficiently. It provides a concise visual summary that allows viewers to grasp the overall layout, spatial relationships, and key features at a glance. Additionally, scene sketching facilitates the creative process by enabling artists and designers to visualize their ideas and concepts.
However, generating a scene sketch is highly challenging, as it requires the ability to understand and depict the visual characteristics of the scene with complex subjects and interactions. (Azadi et al. 2018; Chen, Lai, and Liu 2018; Isola et al. 2017; Li and Wand 2016) often rely on explicit sketch datasets for training. The sketches of them are often simplified and abstract expressions of the original images, with a fixed style or preset. It is difficult to balance visual effect of sketches, producing visual appeal and aesthetics.
Besides, different regions within a scene may have varying levels of importance or prominence. For example, foreground objects or focal points might require more attention to detail and precision in sketching, while background elements may be more loosely represented. Some works can achieve this with flexibility. However, these works focus specifically on the task of object sketching (Muhammad et al. 2018; Vinker et al. 2022b) or portrait sketching (Berger et al. 2013), and often simply use the number of strokes to define the effect of their sketches.
To address the above two issues, we introduce a scene sketching method based on regions with multi-round optimization. We utilize the black parametric Bézier curves as our fundamental shape primitive for strokes of a sketch and optimized them by a pre-trained CLIP-ViT model (Radford et al. 2021; Dosovitskiy et al. 2020) and VGG16 model (Simonyan and Zisserman 2014).
Unlike (Frans, Soros, and Witkowski 2022; Schaldenbrand, Liu, and Oh 2022), optimization is performed for different regions in our method, which helps highlight regions of interest, achieving coarse-to-concrete sketches. An intuitive and succinct learning process can be seen in Figure 1). In each round of optimization, we are in pursuit of full content exploration rather than only the salient guidance. To achieve this, we present an edge-based stroke initialization and utilize a farthest point sampling (FPS) algorithm to uniformly sample strokes in the input image. Figure 5 shows that our sampling method has a more efficient effect and is able to generate sketches in a reasonable manner. To transform a detailed scene into a simplified sketch, one must condense intricate visual elements into fundamental lines, shapes, and tones, all while retaining the scene’s identifiable characteristics. We utilize intermediate layer of CLIP-ViT (Dosovitskiy et al. 2020) to guide the optimization, where encourages the creation of looser sketches that emphasize the scene’s semantics. We introduce a VGG16 (Simonyan and Zisserman 2014) model to promote the visual consistency and similarity between a sketch and an image.
We evaluate our proposed method for various photographs, including people, nature, indoor, animals et al., to showcase the effectiveness of our method. Our main contributions in this work can be summarized as follows:
  • We propose a novel method to convert a scene photograph into a sketch by optimizing it in a region-by-region fashion. A variety of sketching results can be achieved by adjusting regions in our scheme.
  • We apply a farthest point sampling (FPS) algorithm to the input image, evenly sampling positions on region edges as stroke positions, which are wisely utilized to emphasize content information.
  • We introduce two novel loss functions, a CLIP-Based Semantic loss and a VGG-Based Feature loss. These losses improve sketch generation, reflecting a balanced combination of both semantic and geometric features.
@@
Related Work
Photo-Sketch Synthesis. Traditional photo-sketch synthesis methods are mainly based on image processing techniques, such as convolution, filtering, and grayscale, etc., to obtain corresponding sketches style pictures. However, the generated sketch of these methods are of low quality. Researchers investigated approaches that related to the deep learning, considering photo-sketch as a cross domain image-to-image translation problem. (Muhammad et al. 2018) trained a model for abstract sketch generation through reinforcement learning of a stroke removal policy that learns to predict which strokes can be safely removed without affecting recognizability. (Li et al. 2019) proposed a learning-based method to resolve the diversity in the annotation of datasets. Pinz (Kampelmuhler and Pinz 2020) employed a fully convolutional encoder-decoder structure to accomplish a mapping from image space to sketch space.
Yet, recently, CLIPasso (Vinker et al. 2022b) proposed a novel object sketching method that can achieve different levels of abstraction, guided by geometric and semantic simplifications. They defined a sketch as a set of Bézier curves and extracted the salient regions of the input image to define the initial locations of the strokes. In contrast to CLIPasso  (Vinker et al. 2022b), our method is not restricted to objects and can handle the challenging task like scene sketching. Additionally, while they only utilized a single form of optimization, we disentangle optimization into multiple rounds controlling the fidelity of different region in the input image, which allows for a wider range of editing and manipulation.
Vector Graphics. Vector representations are widely used for a variety of sketching tasks and applications, combining with a number of deep learning models including RNN (Ha and Eck 2017), CNNs (Chen et al. 2017), BERT (Lin et al. 2020), Transformers (Bhunia et al. 2020; Ribeiro et al. 2020), GANs (Balasubramanian, Balasubramanian et al. 2019) and reinforcement learning algorithms (Ganin et al. 2018; Mellor et al. 2019; Zhou et al. 2018). While traditional image generation methods operating over vector images require a vector-based dataset, (Li et al. 2020; Mihai and Hare 2021) had shown its possible to manipulate or synthesize vector content by using raster-based loss functions bypassing this limitation (that is, the process of actually drawing the vectors into an image is not part of the learning machinery). Among them, we consider to use the method of (Li et al. 2020), as it can flexibly handle a wide range of curves and strokes, including Bézier curves.
CLIP-Based Image Vectorization.  (Radford et al. 2021) proposed CLIP, which is a neural network trained on 400 million image-text pairs collected from the internet with the objective of creating a joint latent space. Being trained on a wide variety of image domains along with lingual concepts, CLIP models are found to be very useful for a wide range of zero-shot tasks, and have enabled a number of successful methods for drawings, such as CLIPDraw (Frans, Soros, and Witkowski 2022), StyleCLIPDraw (Schaldenbrand, Liu, and Oh 2022), CLIP-CLOP (Mirowski et al. 2022), and CliPascene (Vinker et al. 2022a)(Tian and Ha 2022) employed evolutionary algorithms combined with CLIP, to produce creative abstract concepts represented by colored triangles guided by text or shape. Their results are limited to either fully semantic (using CLIP’s text encoder) or entirely geometric (using L2). CLIPasso (Vinker et al. 2022b) introduced a new loss term and a saliency-guided initialization procedure. Our method also utilizes the CLIP image encoder to guide the process of converting a photograph to a sparse sketch with a combination of features that capture the essence of the input image.
Method
Figure 5: Different Stroke Sampling Method. Top to Bottom: our proposed sampling method (FPS), the sampling method of CLIPasso (Vinker et al. 2022b), random sampling.
Figure 6: The comparisons of scene sketching results. On the right, are three representative sketches produced by our method depicting three levels of abstraction. More comparisons are provided in the supplementary material.
@@
We present a new method to processively convert a given scene photograph into a sketch with multiple rounds of optimization. An overview of our method can be seen in Figure 2. Briefly, given an arbitrary image, our method can recursively learn its different regions by adding optimizable Bézier curves. We define our sketch as some sets of black Bézier curves from different regions placed on a white background. Firstly, we introduce a stroke allocation method to reasonably divide the total number of strokes into different regions. Then we use a proposed stroke sampling to determine stroke locations, which will be converted into our initial strokes (Bézier curves). To improve convergence, we define the order from the global region to other regions to optimize strokes successively.
There are several advantages behind our method. First, it considers the information from different regions of the input image separately, which can help ignore background interference (see Figure 7). Also, the sketch results at the same level of abstraction is greatly increased based on different choices of the regions (see Figure 8). Last, during our optimization process, we also can easily obtain sketches results at different levels of abstraction without changing the total number of strokes (see Figure 1).
Stroke Initialization
@@
Stroke Allocation. Our method is based on multiple rounds of stroke superposition. Thus we first should consider the allocation of strokes in different regions. As a case of fairness, we allocate strokes according to the edge points in each region, based on the information gathered from each region’s edge contents. $EdgeDetector$ represents the edge extractor to gain the number of edge points in the region $R_{i}$. $r$ presents the number of regions. The stroke allocation ratio of region $R_{i}$ is calculated as follows:
@@
Stroke Sampling. As each stroke corresponds to a sample point, the distribution of those points determines the distribution of strokes. For uniform coverage of image information and to reduce the lack of information, we adopt a farthest point sampling (FPS) (the process shown in Figure 4). As a first step, we process the region to get the edges. Next, we use a farthest point sampling to sample strokes, selecting a subset of points from a larger set and maximizing the minimum distance between any two selected points. This process ensures that the selected points are well-spaced and representative of the entire set.
@@
Figure 5 shows that our stroke sampling method contributes significantly to the quality of the final sketch compared to the sampling method of CLIPasso (Vinker et al. 2022b) and random sampling. We further analyze the effect and variability of FPS in the supplementary material.
Loss Function
In previous works, some commonly used loss functions to minimize the error between images and results based on pixel-wise metrics. Although pixel-wise loss is simple yet intuitive, it is not sufficient to measure the distance between sketches and images as a sketch is highly sparse and abstract.
To address this, we leverage a pre-trained CLIP model to guide the training process.
CLIP-Based Semantic Loss. Due to the capabilities of encoding shared information from both sketches and images, we follow the work of (Vinker et al. 2022b) using CLIP model to compute the distance between the embeddings of the sketch $CLIP(Sketch)$ and image $CLIP(Image)$ as:
@@
where $dist(x,y)=1-\frac{x\cdot y}{\|x\|\cdot\|y\|}$ is the cosine distance. However, while (Vinker et al. 2022b) utilize the ResNet-based (He et al. 2016) CLIP model for sketching, we find that the ViT-based (Dosovitskiy et al. 2020) CLIP model provides better coverage of global context. Therefore the loss function is defined as the L2 distance between the activation of ViT-B/32 CLIP model at layers:
@@
VGG-Based Feature Loss. To further improve the similarity between the image and the sketch, we compute the L2 distance between the feature of the sketch and the image based on VGG16 (Simonyan and Zisserman 2014):
@@
For more details of VGG model we used please refer to the supplementary material. The final objective of the optimization is then defined as:
@@
Optimization
For each optimization round, the parameters of stroke are superimposed with the previous round. All parameters are trained for 800 iterations. We define that the first round of optimization is from the global region, and we are able to gain a sketch at a low level of abstraction. Then the sketch will be refined based on the superimposition of strokes from other regions selected by the user. Since the strokes superimposed are the results of stroke sampling from the next region, the sketch in each optimization is flexible.
In Figure 3, we show the process of our optimization and sketch results at different levels of abstraction generated by different optimization round.
Experiments
In the following, we demonstrate the performances of our technique qualitatively and quantitatively, and provide comparisons to state-of-the-art sketching methods. Figure 8 illustrates the impact of the regions. When different regions are chosen, we may get some variations, with the tradeoff that it may ignore the other regions. Further analysis, results, and a user study are provided in the supplementary material.
Ours XoG CLIPasso Photo-Sketching Chan et al. UPDG
LPIPS $\downarrow$ 0.519  0.531  0.566 0.549 0.702 0.628 0.520 0.740
SSIM $\uparrow$ 0.690  0.680  0.654 0.740 0.614 0.634 0.632 0.624
@@
Table 1: Comparison of the LPIPS score and SSIM score. Our scores correspond from left to right to the sketches under the abstraction levels from left to right in Figure 6.
Qualitative Evaluation
Figure 7: Comparison to CLIPasso (Vinker et al. 2022b) without scene decomposition and to CLIPascene (Vinker et al. 2022a) with scene decomposition.
We compare our method with alternative sketching methods for different subjects, including XoG (Winnemöller, Kyprianidis, and Olsen 2012), Photo-Sketching (Li et al. 2019), Chan et al. (Chan, Durand, and Isola 2022) and UPDG (Yi et al. 2020). Note that, none of these sketching approaches can produce sketches with varying abstraction levels, and none can produce sketches in vector format. Due to the fact that our method requires a predefined number of strokes as input, we present three sketches produced by our method depicting three representative levels of abstraction.
The sketches produced by UPDG (Yi et al. 2020) and Chan et al. (Chan, Durand, and Isola 2022) are detailed, closely similar to the edge maps of the input images (like results of XoG (Winnemöller, Kyprianidis, and Olsen 2012)). The geometric structures of these sketches are most similar to the input images shown in the leftmost column of Figure 6. However, some edges and semantic information may be lacking. The sketches produced by Photo-Sketching (Li et al. 2019) are less detailed and difficult to align well with the input scene, which is difficult to identify the main content. Our generated sketches are able to maintain recognizability, underlying structure, and essential visual components of the scene drawn.
In Figure 7, we provide additional comparisons with CLIPasso (Vinker et al. 2022b) and CLIPascene (Vinker et al. 2022a). Both of them can produce sketches with varying abstraction levels and produce sketches in vector format. We generate sketches with the same number of strokes. CLIPasso (Vinker et al. 2022b) aims to portray objects accurately. However, for the scene, it fails to depict the total image in details. This drawback may result from its stroke sampling method based on an attention map.
Then we compared to CLIPascene (Vinker et al. 2022a), for a fair comparison, we use the same decomposition technique to separate the input images into foreground and background. Then we use our method to sketch each part separately before combining them. CLIPascene (Vinker et al. 2022a) captures the whole image, which prevents them from emphasizing important regions of the input image. Our method is able to capture the entire image and display key regions controllably. For example, the dog in the first row.
Figure 8: Comparison of sketches obtained under different regions division at the same number of strokes.

### v2

@@ Abstract
Scene sketching is to convert a scene into a simplified, abstract representation that captures the essential elements and composition of the original scene. It requires a semantic understanding of the scene and consideration of different regions within the scene. Since scenes often contain diverse visual information across various regions, such as foreground objects, background elements, and spatial divisions, dealing with these different regions poses unique difficulties. In this paper, we define a sketch as some sets of Bézier curves because of their smooth and versatile characteristics. We optimize different regions of input scene in multiple rounds. In each optimization round, the strokes sampled from the next region can seamlessly be integrated into the sketch generated in the previous optimization round. We propose an additional stroke initialization method to ensure the integrity of the scene and the convergence of optimization. A novel CLIP-based Semantic Loss and a VGG-based Feature Loss are utilized to guide our multi-round optimization. Extensive experimental results on the quality and quantity of the generated sketches confirm the effectiveness of our method.
Index Terms:
sketch, vector image, optimization, image processing
I Introduction
@@
Scene sketching offers advantage of conveying information quickly and efficiently. It provides a concise visual summary that allows viewers to grasp the overall layout, spatial relationships, and key features at a glance. Additionally, scene sketching facilitates the creative process by enabling artists and designers to visualize their ideas and concepts [1, 2, 3].
However, generating a scene sketch is highly challenging, as it requires the ability to understand and depict the visual characteristics of the scene with complex subjects and interactions. [4, 5, 6, 3] often rely on explicit sketch datasets for training. The sketches of them are often simplified and abstract expressions of the original images, with a fixed style or preset. It is difficult to balance visual effect of sketches, producing visual appeal and aesthetics.
Besides, different regions within a scene may have varying levels of importance or prominence as seen in Fig. 1. For example, foreground objects or focal points might require more attention to detail and precision in sketching, while background elements may be more loosely represented. Some works can achieve this with flexibility. However, these works focus specifically on the task of object sketching [7, 8] or portrait sketching [9, 10], and often simply use the number of strokes to define the effect of their sketches.
To address the above two issues, we introduce a scene sketching method based on regions with multi-round optimization. We utilize the black parametric Bézier curves as our fundamental shape primitive for strokes of a sketch and optimized them by a pre-trained CLIP-ViT model [11, 12] and VGG16 model [13].
Figure 1: Drawings of different scenes by different artists. Notice the significant differences in level of abstraction between different regions of the drawings.
@@
Unlike [14, 15, 16], optimization is performed for different regions in our method, which helps to highlight regions of interest, achieving coarse-to-concrete sketches. An intuitive and succinct learning process can be seen in Fig. 2. In each round of optimization, we are in pursuit of full content exploration rather than only the salient guidance. To achieve this, we present an edge-based stroke initialization and utilize a farthest point sampling (FPS) algorithm to uniformly sample strokes in the input image. Fig. 5 shows that our sampling method has a more efficient effect and is able to generate sketches in a reasonable manner. To transform a detailed scene into a simplified sketch, one must condense intricate visual elements into fundamental lines, shapes, and tones, all while retaining the scene’s identifiable characteristics. We utilize intermediate layer of CLIP-ViT [12] to guide the optimization, where encourages the creation of looser sketches that emphasize the scene’s semantics. We introduce a VGG16 [13] model to promote the visual consistency and similarity between a sketch and an image.
We evaluate our proposed method for various photographs, including people, nature, indoor, animals et al., to showcase the effectiveness of our method. In summary, our contributions encompass four aspects: (1) We propose a novel method to convert a scene photograph into a sketch by optimizing it in a region-by-region fashion. A variety of sketching results can be achieved by adjusting regions in our scheme. (2) We apply a farthest point sampling (FPS) algorithm to the input image, evenly sampling positions on region edges as stroke positions, which are wisely utilized to emphasize content information. (3) We introduce two novel loss functions, a CLIP-based Semantic Loss and a VGG-based Feature Loss. These losses improve the generation of sketches, reflecting a balanced combination of both semantic and geometric features.
II Method
@@
We present a new method to processively convert a given scene photograph into a sketch with multiple rounds of optimization. An overview of our method can be seen in Fig. 2. Briefly, given an arbitrary image, our method can recursively learn its different regions by adding optimiziable Bézier curves. We define our sketch as some sets of black Bézier curves from different regions placed on a white background. Firstly, we introduce a stroke allocation method to reasonably divide the total number of strokes into different regions. Then we use a proposed stroke sampling to determine stroke locations, which will be converted into our initial strokes (Bézier curves). To improve convergence, we define the order from the global region to other regions, optimizing strokes successively.
II-A Stroke Initialization
@@
Stroke Allocation. Our method is based on multiple rounds of stroke superposition. Thus we first should consider the allocation of strokes in different regions. As a case of fairness, we allocate strokes according to the edge points in each region by, based on the edge information gathered from contents of regions by edge detection [17]. $EdgeDetector$ in the following formula represents the edge extractor to gain the number of edge points in the region $R_{i}$. $r$ presents the number of regions. The stroke allocation ratio of region $R_{i}$ is calculated as follows:
@@
Figure 5: Different Stroke Sampling Method. Top to Bottom: FPS sampling method, the sampling method of CLIPasso [8], random sampling.
Stroke Sampling. As each stroke corresponds to a sample point, the distribution of those points determines the distribution of strokes. For uniform coverage of image information and to reduce the lack of information, we adopt a farthest point sampling (FPS) [18] (the process shown in Fig. 3). As a first step, we process the region to get the edges. Next, we use a farthest point sampling to sample strokes, selecting a subset of points from a larger set and maximizing the minimum distance between any two selected points. The process ensures that the selected points are well-spaced and representative of entire set.
@@
Fig. 5 shows that our stroke sampling method contributes significantly to the quality of the final sketch compared to the sampling method of CLIPasso [8] and random sampling. We further analyze the effect and variability of FPS in the supplementary material.
II-B Loss Function
In previous works, some commonly used loss functions to minimize the error between images and results based on pixel-wise metrics. Although pixel-wise loss is simple yet intuitive, it is not sufficient to measure the distance between sketches and images as a sketch is highly sparse and abstract. To address this, we leverage a pre-trained CLIP model to guide the training process.
CLIP-based Semantic Loss. For encoding shared information from both sketches and images, we follow CLIPasso [8] using CLIP model to compute the distance between the embeddings of the sketch $CLIP(Sketch)$ and image $CLIP(Image)$ as:
@@ II-B Loss Function
where $dist(x,y)=1-\frac{x\cdot y}{\|x\|\cdot\|y\|}$ is the cosine distance. And we are also based on the definition of the L2 distance loss function between certain activation layer $l$ in the CLIP model:
@@
VGG-based Feature Loss. To further improve the structural similarity between the image and the sketch, we compute the L2 distance between the feature of them based on VGG16 [13]:
@@
More details about the CLIP model, activation layer and VGG model could be find in the supplementary material. The final objective of the optimization is then defined as:
@@
TABLE I: Comparison of the LPIPS [19] score and SSIM [20] score. The score from left to right correspond to the results of our method, CLIPasso [8] and CLIPascene [14] based on sketch results at low and high abstraction levels.
Ours CLIPasso CLIPascene XoG Photo-Sketching Chan et al. UPDG
LPIPS $\downarrow$ 0.566  0.519 0.672  0.602 0.577  0.525 0.549 0.728 0.520 0.640
SSIM $\uparrow$ 0.654  0.690 0.613  0.622 0.618  0.676 0.740 0.334 0.652 0.649
TABLE II: Comparison of user preference rates. The scores from left to right correspond to the results of our method, CLIPasso [8] and CLIPascene [14] are based on sketching results at low and high abstraction levels.
Ours v.s. CLIPasso Ours v.s. CLIPascene Ours v.s. XoG Ours v.s. Photo-Sketching Ours v.s. Chan et al. Ours v.s. UPDG
Ours 88.7%  96.3% 12.5%  38.6% 41.9% 96.2% 42.5% 40.0%
Others 5.0%  2.5% 10.0%  26.4% 38.4% 3.8% 32.5% 28.8%
Equal 6.3%  1.2% 77.5%  60.0% 19.7% 0.0% 25.0% 32.2%
II-C Optimization
In each optimization round, stroke parameters are overlaid with those from the previous round, with all parameters trained for 800 iterations. We start with a low-level sketch from the global region in the first round of optimization, refining it by incorporating strokes from other user-selected regions. As strokes from subsequent regions are sampled, each optimization round yields a flexible sketch. Fig. 4 illustrates this process, showcasing sketches at varying levels of abstraction across different optimization rounds. It should be noted that multiple rounds, while allowing for careful optimization of regions, can also lead to excessive overlap between regions, affecting the final result.
III EXPERIMENTS AND RESULTS
III-A Implementation
Figure 6: Comparisons to methods that generate pixel-based sketches. On the right, are three representative sketches produced by our method depicting three levels of abstraction. More comparisons are provided in the supplementary material. More comparisons are provided in the supplementary material.
Figure 7: Comparisons to methods that generate vector-based sketches. Representative sketches at high and low level of abstraction are generated.
Optimization Details. We use Adam optimizer with a learning rate set to 1. We evaluate the output sketch every iterations. Evaluation is done by computing the loss without random augmentations. We repeat the optimization process until convergence (when the difference in loss between two successive evaluations is less than 0.00001), this typically takes around 800 iterations. It takes 4 minutes to run 800 iterations on our server NVIDIA RTX 3090 GB, however, after 500 iterations, it is already possible to get a recognizable sketch for most images.
Curves Details. For each curves, we sample positions for their first control points using the FPS algorithm and then randomly sample the next three control points of each Bezier curve within a small radius (0.05) of the first point.
Edge Detection. We use the Canny edge detection [17] on the preprocessed image. It takes the grayscale image as input along with two threshold values: the lower threshold and the upper threshold. We determining the appropriate threshold values as (20, 200) by a trial-and-error process.
III-B Evaluation Metrics
We calculate the Learned Perceptual Image Patch Similarity (LPIPS)  [19] and Structural Similarity Index (SSIM)  [20] to evaluate models. The LPIPS assesses perceptual similarity by computing the L2 distance between deep feature representations extracted from pretrained convolutional neural networks, such as VGG [13]. It measures the perceptual quality and semantic similarity between the synthetic images and the ground-truth images. The SSIM evaluates structural similarity by analyzing luminance, contrast, and structure between image pairs, providing a value between 0 and 1, where 1 indicates identical images.
III-C Comparison to State-of-the-art Methods
We conduct a comparison of our method against various alternative sketching techniques, such as XoG [21], Photo-Sketching [22], Chan et al. [23], and UPDG [24]. It is worth noting that none of these techniques possess the capability to adjust the abstraction level of the sketches or produce vector-based sketches. We also provide comparisons with vector-based methods CLIPasso [8] and CLIPascene [14]. Considering that the sketches generated by XoG [21], Photo-Sketching [22], Chan et al. [23] and UPDG [24] are at different levels of abstraction, for a more comprehensive comparison, we use two different numbers of strokes (128 and 32) as inputs to our method, CLIPasso [8] and CLIPascene [14] to generate sketches at different levels of abstraction. Quantitative results in Table I show our mehod achieves competitive LPIPS and SSIM scores compared to existing methods, indicating high perceptual similarity and structural consistency. While XoG achieves slightly better scores in SSIM (0.740) and LPIPS (0.549), its inability to control abstraction levels limits its flexibility compared to our method.